Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎


Crew 100B Log Book for February 28 2011

Photos of the day

Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Excellent

Experiment Status: All the medical, astrobiological, rover, and food experiments began today with unique results.

The meals today were very delicious, since it was a non-cooking day.

Productivity: We were very innovative today , since our rover engineer and crew scientist continued to work on their experiments .

EVA’s: We completed four successful EVA’s.

Plan for tomorrow: Begin with doing the medical experiment again on all crew members. Observations of the Sun via the radio telescope will also begin. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will occur. Soil samples will again be collected during three EVA’s tomorrow. One will be on foot, and two will use the ATV’s. Tomorrow will be a cooking day.

Commander's Report

Balwant Rai Reporting

Today we began our day with our continuing medical experiments as well as a new examination that looked at every crew member’s dental health.

The crew’s dental health will continue to be monitored by Jasdeep for the remainder of our stay here on Mars. Breakfast was very agreeable and satisfying. The second floor of the HAB was a little messy, so we made it look like a 5-star hotel. Afterwards, some of the crew started the first EVA of the day with creating an obstacle course for the rover, and continuing to get familiar with flying the drone. The rest of our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well.

impromptu introduction to the night’s sky for the entire crew, since it looks like Our crew scientist continued to work through problems with starting observations of sun via the radio telescope. She has been contacting a number of outside astronomers in order to determine why the telescope is not receiving any signals. Tonight, our crew scientist plans to give an it will be a very clear night. Two more EVA’s occurred, where soil samples were taken for the microbiology experiment. Later in the evening one more EVA occurred for a few of our crew members to explore another area of the Martian landscape.

Our resident alien, helped the crew out today by fixing some problems with our Martian home. Today is a non-cooking day, so I am sure our chief and associate chef will whip up something interesting. Tomorrow, we will continue our medical and astrobiological experiments, as well as continue working on getting the rover working well, and potentially getting the radio telescope up and working.

Engineering Report

Matthieu Ansart Reporting
· Kitty (Diesel Generator) (Used? Yes/No): Yes
· Honey (Gas Generator) (Used? Yes/No): No
· Battery Charge Level: Checked by DG
o 10.59 V everstart
o 12.37 V armor plate
o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run, all is good

Fuel Status:
· Diesel (Full/0.75/0.5/0.25/Empty): 0.73
· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.41
· Gasoline Tank (Full/0.75/0.5/0.25/Empty): ?

Notes/Comments: We can't see de gasoline level in the Gasoline Tank but it isn't empty because we fill 2 jerry can

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG
· Honey Oil Quantity (# of Quarts in storage): Leave it to DG
· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench
· Car Oil Quantity (# of Quarts in storage):

Water Status:
· Outside Potable Water Tank Level (inches from bottom): 20
· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full
· Hab Potable Water Tank Level (inches from bottom): 12
· Potable Water Meter Reading:


· Condition of plants in Tank 1: green water
· Condition of plants in Tank 2: green things
· Condition of plants in Tank 3: green things
· Greenhab notes/ comments: Tank 1 is almost full, so soon, Tank 2 and Tank 3 could be filled.

· Telescope Used? (Yes/no) No
Observatory Notes:

· Used (yes/no): No
· Oil Checked (yes/no): No
· Tire Status: Not checked
· Odometer at end of day: 123973
· Notes/Comments on Hab Car:

ATV 1 (Opportunity):
· Used: Yes
· Oil Checked: No
· Fuel Consumed:Replaced
· Tire Status: Filled
· Comments:

ATV 2 (Spirit):
· Used: No
· Oil Checked: No
· Fuel Consumed:
· Tire Status:
· Comments: Still not fixed

ATV 3 (Viking I):
· Used: Yes
· Oil Checked: No
· Fuel Consumed: Replaced
· Tire Status: Filled
· Comments:

ATV 4 (Viking II):
· Used: Yes
· Oil Checked: No
· Fuel Consumed:Replaced
· Tire Status: Filled
· Comments:

Heating and Ventilation:
· Thermostat setting upstairs (Farenheit): 67
· Actual temperature upstairs (Farenheit): 74
· Thermostat setting downstairs (Farenheit): 65
· Actual temperature downstairs (Farenheit): 64

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering
1. Notes/Comments

Science Report (Medical)

Balwant Rai Reporting

Autonomic nervous system variables of EVA Participants of Mars Analogue Mission

Astronauts who have been exposed to microgravity during EVA have some degree of orthostatic hypotension during a postflight stand test, with a reduced mean arterial pressure and excessive cardioacceleration, which indicates a disturbed autonomic nervous system. It is of extreme consequence that the crew of a craft that is to land on Mars be physically fit in order to handle their work tasks and to be prepared for possible emergency situations.

A journey to Mars will take months to complete and as such an appropriate exercise regime must be in place to ensure the Safety and productivity of the crew. To the best my knowledge no evaluation has been made on the autonomic nervous system involved in Mars analogue EVA activity.

Hence this study is planned the effects of EVA on Autonomic nervous system. Autonomic regulation of the cardiovascular system was assessed by measuring the heart rate variability using (via the Zephyr Bioharness, NERV Express 4.2 software) and the spontaneous baroreflex.

All members were selected for this study. The results showed significantly decrease in activity of sympathetic nervous system after EVA. So Marsonaut should be required to participate in physical activity.

Science Report (Oral Hygiene)

Jasdeep Kaur Reporting

Time: 0930-sample collection

Crew Members: Science Team: 6 crew members Balwant Rai, Rachel Dompnier , Quentin Bourges,Chrystal Latham,Matthieu Ansart and Jasdeep Kaur

Site Location: Hab

Objective: To figure out dental findings and other oral findings. PREVENTION IS BETTER THAN CURE

Observation: All crew members were observed for their dental findings and oral findings. All teeth were examined for caries, periodontal infections, gums, saliva, tongue, mucosa, TMJ problems. Mouth opening was noted for all members.

Results: As with respect to the principles to be followed on Mars, crew members cannot use non-biodegradable products, so toothpastes, floss, mouthwashes cannot be used and crew members have to follow the rules to brush the teeth with baking soda. This needs very good maintainance of teeth. Also, few members experienced mild pain due to impacted teeth, so they were guided to follow some exercises and saline rinses. All crew members have healthy gingiva, excellent periodontal status and have no acute infections or any pathology.

Brushing Technique: Brushing technique was explained to all crew members for proper maintainance of oral health.

EVA 4 Report

Rachel Dompnier Reporting

EVA #4, 10:30-11:20

Crew Members: Matthieu Ansart, Quentin Bourgès, Rachel Dompnier

Site Location: Behind the Hab

Transit Mode: Foot

Objectives: Getting acclimated to controlling the drone and to test its accuracy outside.

Results: Controlling the drone is not ever easy because it moves fast and sometimes not obeys the orders from the operator.

Lessons Learned: The drone is not so accurate, especially when it is windy. We need to be more aware of its environment.

EVA 5 Report

Jasdeep Kaur Reporting

EVA #5, 12:15-13:00

Crew Members: Matthieu Ansart, Crystal Latham, Jasdeep Kaur, Rachel Dompnier , Quentin Bourges

Site Location: N 38.40651, W 110.78990

Transit Mode: Foot

Objectives: GeoMission : to collect soil samples with designed scientific protocols to look for confirmation of microbes and extremophiles.

Results: Sample was collected in a direction perpendicular to Hab in N-W direction. This area exhibited characteristics of cracks, dried impressions as well as ground fissures in an area that had a very peculiar color.

Lessons Learned: All sterilization techniques and protocols to be followed

EVA 6 Report

Jasdeep kaur reporting

EVA #6, 14:15-15:00

Crew Members: Balwant Rai, Rachel Dompnier , Crystal Lantham

Site Location: N 38.40519, W 110.79529

Elevation: 1398mts

Transit Mode: ATVs

Objectives: GeoMission to collect soil samples with designed scientific protocols to look for confirmation of microbes and extremophiles

Results: Collected a sample of greenish-yellow rock/ and soil that was in an area that exhibited characteristics of splinters and fissures

Lessons Learned: Be sure to follow all protocols when collecting samples in a sterilized way without any contamination.

EVA 7 Report

Matthieu Ansart reporting
EVA #7, 17:05-17:45
Crew Members: Quentin Bourgès, Matthieu Ansart
Transit Mode: ATVs
Objectives : Recognition : to recognize the area around the road at the East and at the West of the hab for about 2 kilometers.
Results : A better knowledge of the area.
Lessons Learned: Radio communication with mission support in the Hab isn't working far from the station. It could be dangerous for our own security. So we always have to be two or more (it's better) for an EVA by ATV in order to prevent bad situations (accidents...).

Chef's Report

Matthieu Ansart Reporting
1a) Today was a non-cooking day
1b) Was today a special day celebrated at the main meal with special food or activity?
2) List the foods served at today's main meal (usually the evening meal), giving full names of each.
Mushroom pilaf with vegetables and chicken noodle soup.
3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)
Mushroom pilaf with vegetables (10 percent of initial quantity)
4) Recipes for anything more complicated than following package directions exactly, or rehydrating.
5) Lessons learned (if any)
6) Comments/questions for Kim and Jean, the food study investigators

Journalist Report

Crystal Latham Reporting

It is crew 100B’s second full day on Mars and it was very exciting. Our medical tests in the morning have been going well, and a new examination was given today. Each crew member had their teeth checked for cavities and overall health.

A few crew members participated in the first EVA of the day, where an obstacle course was made for the rover. It was a very beautiful day here on Mars, so the drone got to stretch its propellers a bit as well. Our crew astronomer has been trying to get the radio telescope operational, with no avail so far. With some help from some astronomers back on Earth, it should soon be ready for observations of the Sun by the end of the Martian week.

A number of EVA’s occurred today involving all crew members. The first had crew members walking to a unique location to collect soil samples. Another EVA occurred later in the afternoon to collect more soil samples. On this EVA, crew members trekked through the rough Martian landscape on the ATV’s. It was rough going in some places, but the final view of the Martian landscape, as well as an interesting soil sample, made the trip very worthwhile.

An alien also made a visit here at MDRS to help with some Hab maintenance issues. Our main meal is currently being made, and it will probably involve ketchup. Ketchup seems to be the condiment of choice, no matter what is being eaten. Mixing oatmeal, cereal, milk, and dried fruit also seems to be the breakfast of Marsonauts. Our meal choices are limited, but the crew has been making due.

Tonight, we plan to get a little bit of R&R as well as a good night’s sleep. If the weather continues to cooperate there may be a stargazing session in order. MDRS signing off, hoping to see some more aliens soon!